OpenStudy (josee):

lim of x to 1 of (sqrt(2-x)-1)/(sqrt(5-x)-2)

OpenStudy (amistre64):

youwant an answer? or steps?

OpenStudy (josee):

stepsss

OpenStudy (amistre64):

play with conjugates when dealing with sqrts \[\lim_{x\ to\ 1} \frac{\sqrt{2-x}-1}{\sqrt{5-x}-2}\]

OpenStudy (amistre64):

\[\lim_{x\ to\ 1} \frac{\sqrt{2-x}-1}{\sqrt{5-x}-2}*\frac{\sqrt{5-x}+2}{\sqrt{5-x}+2}\] \[\lim_{x\ to\ 1} \frac{\sqrt{(2-x)(5-x)}-\sqrt{5-x}+2\sqrt{2-x}-2}{5-x-4}\]

OpenStudy (amistre64):

\[\lim_{x\ to\ 1} \frac{\sqrt{(2-x)(5-x)}-\sqrt{5-x}+2\sqrt{2-x}-2}{1-x}\] doesnt always work out nice but most of the time it does :)

OpenStudy (amistre64):

what methods you allowed to use? LHopital?

OpenStudy (josee):

no derivatives

OpenStudy (josee):

:/

OpenStudy (nikvist):

\[\lim_{x\rightarrow 1}\frac{\sqrt{2-x}-1}{\sqrt{5-x}-2}= \lim_{y\rightarrow 0}\frac{\sqrt{1-y}-1}{\sqrt{4-y}-2}=\frac{1}{2}\,\lim_{y\rightarrow 0}\frac{\sqrt{1-y}-1}{\sqrt{1-y/4}-1}=\] \[=\frac{1}{2}\,\lim_{y\rightarrow 0}\frac{1-y/2-1}{1-y/8-1}=\frac{1}{2}\cdot 4=2\]